OpenStudy (anonymous):

Solve the following system. (Hint: Let x = 1/u and y=1/v. Rewrite the system in terms of x and y and solve. Then use x and y to find u and v.) 6/u + 3/v =2 2/u - 9/v = 4

OpenStudy (anonymous):

x = 1/2 and y = -1/3

OpenStudy (anonymous):

@Gowthaman Please do not just give the answers...